OpenStudy (anonymous):

what is slope intercept bof the line (-1 ,1) and( 0,3)

OpenStudy (anonymous):

k=(3-1)/(0-(-1))=2

OpenStudy (anonymous):

y-y1/y2-y1 = x-x1/x2-x1 (-1,1)&(0,3) y-1/3-1 = x+1/0+1 y/3 = x+1 y=3x+3 the slope is 3 then (y=mx+c)

OpenStudy (anonymous):

thank you